40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 30 06:25:20 2011, kiwamu, Update, LSC, Calbiration of BS, ITMs and PRM actuators actuators.png
    Reply  Sat Oct 8 00:44:42 2011, kiwamu, Update, LSC, calibration of SRM actuator actuators.png
Message ID: 5637     Entry time: Sat Oct 8 00:44:42 2011     In reply to: 5583
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: calibration of SRM actuator 

The AC response of the SRM actuator has been calibrated.

 actuators.png
(Summary of the calibration results)
     BS = 2.190e-08 / f2     [m/counts]
     ITMX  = 4.913e-09 / f2   [m/counts]
     ITMY  = 4.832e-09 / f2   [m/counts]
     PRM   = 2.022e-08 / f2  [m/counts]
     SRM   = 2.477e-08 / f2  [m/counts]    ( NEW ! ) 
 
(Measurement)
The same technique as I reported some times ago (#4721) were used.
The Signal-Recycled ITMY was locked for measuring the actuator response.
Since the ITMY actuator had been already calibrated, first the sensor was calibrated into [counts/m] by exciting the ITMY actuator and then calibrated the SRM actuator with swept sine measurement.
 
 - - notes to myself
   SRCL GAIN = 2.2
   Sensor = REFL11_I
   Demod. phase = 40 deg
   Resonant condition = Carrier resonant
   Gain in WF = 0 dB

Quote from #5583
The AC responses of the BS, ITMs and PRM actuators have been calibrated.

 

ELOG V3.1.3-